OpenStudy (anonymous):

Mai bought a television for $325. The tax rate is 8%. Which proportion can be used to find the amount of tax Mai will pay?

OpenStudy (anonymous):

8/100 = x/35?

OpenStudy (phi):

I would write 8% as a fraction and tax/total_price as the other fraction (we could call tax "x", for unknown)

OpenStudy (anonymous):

\[\frac{ 8 }{ 100 } = \frac{ x }{ 35 }\]

OpenStudy (phi):

what is 35 ?

OpenStudy (anonymous):

err 325 not 35

OpenStudy (phi):

ok, that makes more sense

OpenStudy (anonymous):

x = 26

OpenStudy (phi):

yes, the tax is 26

OpenStudy (anonymous):

so the way to cross multiply is the way i did it? or is it a different way?

OpenStudy (phi):

people learn "cross multiply" but the easiest way is this \[ \frac{ 8 }{ 100 } \cdot \cdot 35= \frac{ x }{ 35 } \cdot 35 \] the idea is if you multiply the right side by 35, the 35/35 turns into 1 but to keep things equal, multiply the left side by 35 also

OpenStudy (anonymous):

like \[\frac{ 8 }{ 325 } = \frac{ x }{ 100 }~ or~ \frac{ 8 } { 325 } = \frac{ 100 }{ x }\]

OpenStudy (anonymous):

oh ok and it is 325 not 35.

OpenStudy (phi):

so \[ x = \frac{8}{100} \cdot 35 \]

OpenStudy (anonymous):

ok thanks for the help.

OpenStudy (anonymous):

fan / medal

OpenStudy (phi):

\[ \frac{ 8 }{ 325 } = \frac{ x }{ 100 }~ or~ \frac{ 8 } { 325 } = \frac{ 100 }{ x } \] these don't make a lot of sense to me I would think: 8% means 8 out of 100 , or more exactly 8 dollars of tax divided by 100 dollars of cost so I would write a ratio of tax divided by cost 8/100 and x (unknown tax)/ 325 (known cost)

OpenStudy (anonymous):

oh ok :) thanks again @phi